### v1

# Efficient and Reliable Hitting-Set Computations for the Implicit Hitting Set Approach

Abstract
The implicit hitting set (IHS) approach offers a general framework for solving computationally hard combinatorial optimization problems declaratively. IHS iterates between a decision oracle used for extracting sources of inconsistency and an optimizer for computing so-called hitting sets (HSs) over the accumulated sources of inconsistency. While the decision oracle is language-specific, the optimizers is usually instantiated through integer programming. We explore alternative algorithmic techniques for hitting set optimization based on different ways of employing pseudo-Boolean (PB) reasoning as well as stochastic local search. We extensively evaluate the practical feasibility of the alternatives in particular in the context of pseudo-Boolean (0–1 IP) optimization as one of the most recent instantiations of IHS. Highlighting a trade-off between efficiency and reliability, while a commercial IP solver turns out to remain the most effective way to instantiate HS computations, it can cause correctness issues due to numerical instability; in fact, we show that exact HS computations instantiated via PB reasoning can be made competitive with a numerically exact IP solver. Furthermore, the use of PB reasoning as a basis for HS computations allows for obtaining certificates for the correctness of IHS computations, generally applicable to any IHS instantiation in which reasoning in the declarative language at hand can be captured in the PB-based proof format we employ.
0.1 Introduction
The implicit hitting set (IHS) approach [LS08] offers a general framework for solving computationally hard combinatorial optimization problems declaratively. Successful practical instantiations of IHS include solvers for maximum satisfiability [DB11, DB13a, DB13b, SBJ16], finite-domain constraint optimization [DB13c], answer set programming [SDAJ18], and pseudo-Boolean optimization [SBJ21, SBJ22], as well as other computationally hard problems such as computing unsatisfiable subsets of formulas [IPLM15, GBG23] and quantified Boolean formulas [NMBJ22], propositional abduction [IMM16, SWJ16], and computing optimal causal graphs [HSJ17].
Constituting a hybrid approach, the IHS framework consists of two main components: one for extracting sources of inconsistency in the declarative specification at hand (referred to as core extraction, with cores being constraints ruling out sources of inconsistency), and one for computing so-called hitting sets over sources of inconsistency (the hitting set optimizer). IHS iterates between these two components, accumulating cores and computing hitting sets over the so-far accumulated cores. The hitting set obtained from the optimizer is used to focus the core extraction step (which reasons only on the original declarative specification) on finding sources of inconsistency that eliminate the current best candidate solution. In this way, the instance solved by the hitting set optimizer keeps on growing as more and more cores are extracted. The core extractor is implemented using a decision oracles specific to the declarative language at hand. In contrast, the optimizer is quite standardly instantiated through integer programming. Indeed, despite the fact that there been has significant early work put into investigating effective ways of realizing the hitting set component in IHS especially in the context of MaxSAT [DB11, DB13a], the use of IP solvers remains largely the main mechanism used in modern IHS implementations for efficiency and for computing optimal solutions Further improvements in the hitting set component are generally applicable to IHS implementations.
In this work we explore alternative algorithmic techniques for hitting set computations within IHS based, with the key aim of realizing effective and trustworthy hitting set computations. In terms of algorithmic alternatives to using IP solvers, we explore different ways of employing pseudo-Boolean (PB) reasoning [RM09] as well as stochastic local search [CCL23a, CCL+23b] and their combinations together with IP solving. In terms of trustworthiness, it should be noted that the hitting set optimizers is in charge of ensuring that the objective function at hand is considered exactly. While IP solvers remain today the de facto approach to hitting set optimization within IHS, they are known to sometimes suffer from numerical instability, resulting in erroneous solutions [CKSW13]. While numerically exact implementations of IP solvers exist [EG21, CKSW13], their performance remains today far from the efficiency of IP solvers in their default settings without guarantees on exactness.
The alternative HS computation approaches are applicable in general in different instantiations of the IHS approach, and—being based on exact reasoning—offer correctness guarantees. We extensively evaluate the practical feasibility of the alternative algorithmic approaches to hitting set optimization we propose in the context of pseudo-Boolean (0–1 IP) optimization as one of the most recent instantiations of IHS. Highlighting a trade-off between efficiency and reliability, while a commercial IP solver turns out to remain the most effective way to instantiate HS computations in this setting, due to potential numerical instability using an off-the-shelf IP solver as-is is not generally exact; in fact, we show that exact HS computations instantiated via PB reasoning can be made competitive with a numerically-exact IP solver. Furthermore, the use of PB reasoning as a basis for HS computations allows for obtaining certificates of correctness of IHS computations. As such, one of the contributions of this paper is the development of the first certifying instantiation of a state-of-the-art IHS approach. To achieve this, we build on the VeriPB proof format [BGMN23], which has recently been shown to be applicable to a rich variety of MaxSAT algorithms [VDB22, BBN+23, BBN+24], with IHS being the most important missing paradigm. Our results on certifying IHS are applicable to any IHS instantiation where the reason of the decision oracle can be captured with PB-based reasoning, and as such includes e.g. MaxSAT as well.
0.2 Preliminaries
0.2.1 Pseudo-Boolean Optimization
A literal $\ell$ is a Boolean variable $x$ or its negation $\overline{x}=1-x$, where variables take values 0 (false) or $1$ (true). A pseudo-Boolean (PB) constraint $C$ is a 011 linear inequality ${\textstyle\sum}_{i}a_{i}\ell_{i}\geq A$, where $a_{i}$ and $A$ are integers. Without loss of generality, we often assume PB constraints to be in normal form, meaning that the literals $\ell_{i}$ are over distinct variables, each coefficient $a_{i}$ is positive and the degree (of falsity) $A$ is non-negative. A pseudo-Boolean formula $F$ is a conjunction $\bigwedge_{j}C_{j}$ of PB constraints. When convenient, we view $F$ as a set of constraints. An objective $O$ is an expression $\sum_{i}w_{i}\ell_{i}+lb$ where the coefficients $w_{i}$ and the constant $lb$ are integers ($lb$ stands for “lower bound”: when the $w_{i}$ are all possible, this constant term is indeed a lower bound on the cost).
An assignment $\alpha$ maps variables to either 0 or $1$. The assignment $\alpha$ satisfies a (normalized) PB constraint $C={\textstyle\sum}_{i}a_{i}\ell_{i}\geq A$ if ${\textstyle\sum}_{i}a_{i}\alpha(\ell_{i})\geq A$ and is a solution to a formula $F$ if it satisfies all of its constraints. The cost $\alpha(O)$ of an assignment $\alpha$ under an objective $O=\sum_{i}w_{i}\ell_{i}+lb$ is $\sum_{i}w_{i}\alpha(\ell_{i})+lb$. An instance of the pseudo-Boolean optimization problem is a tuple $(F,O)$ with $F$ a PB formula and $O$ an objective to minimize. The solutions of $(F,O)$ are the solutions of $F$. A solution $\alpha$ is optimal if $\alpha(O)\leq\beta(O)$ for each solution $\beta$ of $(F,O)$. The optimal cost of $(F,O)$ is $\alpha(O)$ for an optimal solution of $(F,O)$.
Following [SBJ21], a constraint $C$ is an core of $(F,O)$ if the following conditions hold: (i) the literals in $C$ are objective literals or their negations and (ii) all assignments $\alpha$ that satisfy $F$ also satisfy $C$ (denoted by $F\models C$). In other words, a core is an implied constraint that only mentions variables occurring in the objective. [^1] This definition of a core comes from the first papers on IHS for PBO [SBJ21]. It should be noted that in the context of e.g. Boolean satisfiability (SAT), an “unsatisfiable core” is standardly defined as a subset of clauses which are unsatisfiable. When phrasing optimization problems such as PBO or MaxSAT in terms of optimizing a cost function subject to (hard) constraints, cores are more conveniently defined as entailed constraints over objective variables. This is particularly evident in the context of PBO where cores can be more generally PB constraints than mere clausal at-least-one cardinality constraints as in MaxSAT, as well as IHS as the cores under our definition are the constraints which are accumulated in the “hitting set” component of IHS.
The following proposition forms the basis for how the IHS approach to PBO uses cores to compute optimal solutions.
Proposition 1 (e.g., [SBJ21]).
Let $(F,O)$ be a PBO instance, $\alpha_{\mathit{best}}$ an optimal solution to $(F,O)$, $\mathcal{K}$ a set of cores of $(F,O)$, and $\gamma$ an optimal solution to $(\mathcal{K},O)$. Then $\gamma(O)\leq\alpha_{\mathit{best}}(O)$.
In words, Proposition 1 states that the cost of the optimal (minimum-cost) solutions to any set of cores is a lower bound on the optimal cost of the instance.
0.2.2 The Implicit Hitting Set Approach
1 $\textnormal{{PBO-IHS}}(F,O)$
 Input: A PBO instance $(F,O)$
 Output: An optimal solution $\alpha_{\mathit{best}}$
2 $(\alpha_{\mathit{best}},\mbox{{sat?}})\leftarrow\textnormal{{PB-Solve}}(F)$;
3 if not sat? then
4    return “no feasible solutions”;
5 $ub\leftarrow\alpha_{\mathit{best}}(O)$; $lb\leftarrow$-$\infty$; $\mathcal{K}\leftarrow\textnormal{{seed}}(F)$;
6 while true do
7    $(\gamma,lb)\leftarrow\textnormal{{Solve-HS}}(\mathcal{K},O,lb,ub)$;
8    if $ub=lb$ then  break;
9    $(\mathcal{N},\alpha)\leftarrow\textnormal{{Extract-Cores}}(\gamma,F,O)$;
10    if $\alpha(O)<ub$ then  $ub\leftarrow\alpha(O)$; $\alpha_{\mathit{best}}\leftarrow\alpha$;
11    if $ub=lb$ then  break;
12    $\mathcal{K}\leftarrow\mathcal{K}\;\cup\;\mathcal{N}$;
13 return $\alpha_{\mathit{best}}$;
Algorithm 1 IHS in the context of PBO.
Algorithm 1 details PBO-IHS, the implicit hitting set algorithm for pseudo-Boolean optimization [SBJ21, SBJ22]. Given a PBO instance $(F,O)$, PBO-IHS begins by checking the existence of solutions of $(F,O)$ by invoking PB decision procedure PB-Solve on $F$ (Line 1). The call returns an indicator sat? for satisfiability and a solution $\alpha_{\mathit{best}}$ in the positive case. Given $\alpha_{\mathit{best}}$, an upper bound $ub$ and a lower bound $lb$ on the optimal cost of the instance are initialized to $\alpha_{\mathit{best}}(O)$ and -$\infty$, respectively, on Line 1. During the search, $\alpha_{\mathit{best}}$ will always contain the currently best-known solution for which $\alpha_{\mathit{best}}(O)=ub$. A set $\mathcal{K}$ of cores is initialized by seeding to be the set of all constraints in $F$ that only mention objective literals.
The main search loop (Lines 1-1) iterates until $lb=ub$, at which point $\alpha_{\mathit{best}}$ is known to be optimal. Each iteration begins with the hitting set [^2] While the term “hitting set” is still standardly used widely in the literature on IHS, is should be noted that the constraints the hitting set component deals with are not at-least-one cardinality constraints hitting as in the classical hitting set problem. In particular, both seeding and the generality of PBO cores results in including more general PB constraints in the “hitting set computations” than mere at-least-one cardinality constraints. component of IHS, i.e., the procedure Solve-HS computing a solution $\gamma$ of the instance $(\mathcal{K},O)$ consisting of the cores found so far and the objective (Line 1). The Solve-HS procedure is the main study of interest in the current paper; we will detail various instantiations of Solve-HS in Section 0.3. In addition to $\gamma$, each call to Solve-HS returns a lower bound $lb$ on the optimal cost of $(F,O)$. As detailed in Section 0.3, the instantiations we consider will return either the input parameter $lb$ unchanged, or the cost $\gamma(O)$ when $\gamma$ is known to be an optimal solution of $(\mathcal{K},O)$ (recall Proposition 1). In order to guarantee the termination of Algorithm 1 with an optimal solution to $(F,O)$ we assume that on each iteration $j$ there exists a $k\geq 0$ s.t. Solve-HS computes an optimal solution of $(\mathcal{K},O)$ (and thus updates $lb$) on iteration $j+k$. For more details and a formal proof of correctness of Algorithm 1, we refer the reader to [BHJS17].
Next, the procedure Extract-Cores is used to obtain new cores. This is done by repeated invocations of a decision procedure for PB constraints of $F$ under a monotonically decreasing set of assumptions $\mathcal{A}$, initialized to the solution $\gamma$. If $F$ is unsatisfiable under $\mathcal{A}$ the decision oracle uses standard conflict analysis methods to extract a new core $C$ of $F$ that contradicts $\mathcal{A}$ and adds it to the set $\mathcal{N}$ of new cores to be returned by Extract-Cores and added to $\mathcal{K}$ in Line 1 of Algorithm 1. Before the next call to the decision procedure, the set $\mathcal{A}$ is relaxed by the so-called weight-aware core extraction [BJ17] heuristic that—informally speaking—removes at least one of the literals in $C$ from the assumptions, thus preventing $C$ from being rediscovered in later iterations. The Extract-Cores procedure terminates when the decision procedure provides a solution $\alpha$ of $F\land\bigwedge_{\ell\in\mathcal{A}}\ell$. While there are no guarantees it will, the solution $\alpha$ can witness an improved upper bound on the cost of $(F,O)$, which is why the upper bound is updated and termination checked on Lines 1 and 1 of Algorithm 1.
0.3 Alternatives Approaches to Hitting Set Computations
We now turn to investigating various alternative instantiations of the hitting set component (i.e., the Solve-HS procedure) for computing optimal solutions to $(\mathcal{K},O)$. Currently efficient instantiations of Solve-HS relatively standardly make use of a commercial IP solver and floating-point computations. This holds true for the current state-of-the-art PBO-IHS solver, implementing IHS for PBO, as well as state-of-the-art IHS implementations for other declarative paradigms [SBJ21, SBJ22, DB13a, DB13c, SDAJ18, GBG23, SWJ16, HSJ17]. While numerically-stable IP solvers are available today, most IP solvers in their default settings employ floating-point computations for practical efficiency, essentially giving up guarantees on finding optimal solutions. This can and has caused intrinsic problems in terms of correctness of IHS. Towards even higher levels of correctness guarantees, verifiable proofs of optimality would be ideal. With these motivations, we explore alternative instantiations of Solve-HS based on recent developments in conflict-driven pseudo-Boolean solvers [DGD+21, DGN21] that allow producing certificates of correctness of the computations of Algorithm 1, as well as various ways of combining them within the hitting set component with IP solving and SLS.
0.3.1 A General View on Hitting Set Solving
1 $\textnormal{{Solve-HS}}(\mathcal{K},O,lb,ub)$
2 if use-sls () then
3    $\gamma\leftarrow\textnormal{{Core-SLS}}(\mathcal{K},O)$
4    if $\gamma(O)<ub$  then
5       return $(\gamma,lb)$
6 $\text{opt?}\leftarrow\textnormal{{optimal-sol}}()$
7 $\gamma\leftarrow\textnormal{{Optimize}}(\mathcal{K},O,ub,\text{opt?})$;
8 if $\gamma(O)=ub\text{ {or} opt?}$  then return $(\gamma,\gamma(O))$ ;
9 else return $(\gamma,lb)$;
10
Algorithm 2 General view on the hitting set component, i.e., computing solutions to the cores extracted so far.
Algorithm 2 details a generalized abstraction of the Solve-HS procedure (the hitting set component), allowing for different ways of combining IP solving, conflict-driven pseudo-Boolean solving, and stochastic local search. Assume PBO-IHS (Algorithm 1) is tasked with computing an optimal solution to an instance $(F,O)$. It will invoke Solve-HS with the set $\mathcal{K}$ of cores extracted so far, the objective $O$, and the current upper and lower bound $(ub,lb)$ on the optimal cost of $(F,O)$.
First Solve-HS invokes use-sls, a user-specified heuristic for determining whether SLS should be invoked on Line 2. If the solution $\gamma$ that SLS returns has a cost lower than the current upper bound $ub$, Solve-HS terminates on Line 2. Otherwise, the subroutine Optimize is invoked on Line 2 and tasked to either compute a solution $\gamma$ of $(\mathcal{K},O)$ that has $\gamma(O)<ub$, or (if no such solutions exist) an optimal solution $\gamma$ of $(\mathcal{K},O)$. Since the current best solution $\alpha_{\mathit{best}}$ of $(F,O)$ maintained by PBO-IHS has $\alpha_{\mathit{best}}(O)=ub$ and is also a solution of $(\mathcal{K},O)$, an alternative view on Optimize is that of finding a solution of $(\mathcal{K},O)$ of cost lower than $ub$, or determining that one does not exist. As an additional heuristic, we allow forcing Optimize to compute an optimal solution to $(\mathcal{K},O)$, essentially forcing an improvement to the lower bound. More specifically, the user-specified heuristic optimal-sol invoked on Line 2 returns true if Optimize is only allowed to terminate with an optimal solution to $(\mathcal{K},O)$.
Before terminating, Solve-HS checks whether the lower bound can be improved to $\gamma(O)$ on Line 2. This can happen if either Optimize was forced to compute an optimal solution or if it was unable to find a solution that has a cost lower than $ub$. Otherwise, the procedure terminates and returns $\gamma$ and the input bound $lb$ on Line 2.
Next, we detail the different parts of Solve-HS the different instantiations of Optimize that we consider as well as integration of local search for potential practical runtime improvements. Our general view on computing hitting sets allows for combining these instantiations in various different ways.
Optimize using an IP Solver.
The arguably most commonly used instantiation of Optimize invokes a commercial MIP solver on $(\mathcal{K},O)$. We consider two variants of IP solving for instantiating Optimize. The first is IP solving techniques that make use of fast floating-point arithmetic. These are arguably among the most efficient methods of computing solutions $(\mathcal{K},O)$, and most state-of-the-art IHS approaches instantiate Solve-HS with a floating-point IP solver. One disadvantage of such solvers, however, is that there have been repeated reports of such solvers outputting faulty solutions [CKSW13], preventing their use in applications that require trust in the results. Thus, we also consider exact IP solving, i.e., IP solving methods that employ exact arithmetic to circumvent issues related to floating point precision and rounding errors [CKSW13, EG23]. This increased reliability of the results does, however, come at a cost in terms of computational efficiency. Recent computational studies [EG23] report slowdowns of a factor $8.1$ when comparing a state-of-the-art exact MIP solver to its floating point counterpart.
Optimize using Conflict-Driven PB Optimization.
Complementing the Branch & Cut search commonly employed by IP solvers, we also consider harnessing recent developments in native pseudo-Boolean optimization procedures that reduce the problem of computing an optimal solution of $(\mathcal{K},O)$ into a sequence of decision problems, which are tackled with a decision procedure similar to the one employed within Algorithm 1 to extract cores [DGD+21, DGN21]. More specifically, we consider three different variants of pseudo-Boolean optimization algorithms for computing an optimal solution to $(\mathcal{K},O)$: solution-improving search (SIS), core-guided search (CG), and core-boosted search (CB).
A SIS-based approach maintains an upper bound $ub_{c}$ on the optimal cost of $(\mathcal{K},O)$, In the context of Solve-HS, $ub_{c}$ is instantiated to the input upper bound $ub$. On each iteration, the decision procedure is invoked on $\mathcal{K}\land(O<ub_{c})$, i.e., the cores in $\mathcal{K}$ and a solution improving constraint $(O<ub_{c})$. If a solution exists, the returned solution $\gamma$ will have $\gamma(O)<ub_{c}$, so the upper bound is updated, and the process repeated. Otherwise (if the decision procedure reports unsatisfiable), the latest solution found is determined to be optimal for $(\mathcal{K},O)$. In addition to the basic SIS approach just described, we consider a variant that we call reified SIS. An incremental SIS approach adds the reified solution-improving constraint $r\Rightarrow(\gamma(O)<ub_{c})$, i.e., the linear inequality $M\overline{r}+\gamma(O)<ub_{c}$ where $r$ is a fresh variable and $M$ is a large enough constant, and invokes the decision procedure while assuming $r=1$. The benefit of relaxing the solution improving constraint is that the decision procedure instantiation can be incrementally maintained between iterations of Solve-HS.
A core-guided approach [DGD+21] maintains a reformulated objective $O^{R}$, initialized to $O$, and a set $\mathcal{C}$ of additional constraints, initialized to $\emptyset$. On each iteration, the decision procedure is queried for a solution of $\mathcal{K}\land\mathcal{C}$ while assuming all literals in $O^{R}$ to 0. More informally, the query asks for a solution $\gamma$ of $\mathcal{K}\land\mathcal{C}$ that has $\gamma(O^{R})=0$. If a solution exists, the obtained solution is an optimal solution of $(\mathcal{K},O)$ and the procedure terminates. Otherwise, the call returns a new core falsified by the assumptions. Next, the core-guided approach introduces new constraints to $\mathcal{C}$ and reformulates $O^{R}$ in a manner that resolves the inconsistency represented by the core in a cost-minimal way. Informally speaking, the invariant maintained by a core-guided search is that optimal solutions of $(\mathcal{K}\land\mathcal{C},O^{R})$ are optimal solutions of $(\mathcal{K},O)$. Since the set $\mathcal{K}$ of cores that we invoke core-guided search on monotonically increases during search, the state of the core-guided solver is maintained between invocations. While such incremental invocations of core-guided search prevent fixing any variables based on the bounds on the optimal cost of $(\mathcal{K},O)$, variables are still fixed based on the bounds of the full instance $(F,O)$. For a concrete example, we fix the value of an objective literal $\ell$ that has a constant larger than $ub$ to 0 since any solution that assigns $\ell=1$ would incur a cost larger than ubub
Finally, core-boosted search [BDS19] is a hybrid of core-guided and solution-improving search that starts with core-guided search for a user-specified number of iterations. Then it invokes solution-improving search on the final reformulated instance $(\mathcal{K}\land\mathcal{C},O^{R})$ obtained by core-guided search.
It is important to note that SIS, CG and CB can each provide intermediate solutions that allow terminating search as soon as a solution of cost lower than $ub$ is found. Specifically, the instantiation of core-guided search that we use makes use of the so-called stratification heuristic [ABGL13] that results in it obtaining intermediate solutions, even if pure core-guided search only ever finds an optimal solution.
0.3.2 Integrating Stochastic Local Search
Stochastic local search (SLS) is a lightweight heuristic approach to pseudo-Boolean optimization that quickly finds low-cost solutions. In contrast to the approaches described in Section 0.3.1, SLS alone can not give any guarantee on optimality of a solution and cannot therefore establish a lower bound on its own. Hence an IHS approach with an SLS integration in the hitting set component cannot terminate without an instantiation of Optimize that can check whether the solution found by SLS is optimal.
Our SLS component of hitting set computations (the Core-SLS procedure in Algorithm 2) is designed to achieve two desirable properties of SLS search in the context of Solve-HS. First, the search should incur as little overhead as possible to justify replacing calls to Optimize that can result in improved bounds on the optimal cost. Second, the solutions found by SLS should, intuitively, be as helpful as those found by Optimize in guiding IHS search toward an optimal solution of the full instance.
Our instantiation of Core-SLS is built on recent advances in SLS for pseudo-Boolean optimization [CCL+23b] and adapts the NuPBO approach to the hitting set context. NuPBO starts from a complete assignment that is not necessarily a solution. In each iteration, the search flips the value of one variable in the hopes of eventually finding a solution. The choice of variable is based on a score that considers three aspects: (i) the contribution of the current variable value to satisfying the cores, (ii) how much flipping the variable value would contribute towards satisfying the cores, and (iii) how flipping the variable value would impact the cost of the assignment.
In more detail, we design Core-SLS for incremental calls, to run in conjunction with another instantiation of Optimize, and to produce good solutions for core extraction. Adapting the SLS approach to incremental calls is straightforward, as we can keep the state from the previous call and simply add the new constraints representing additional cores. Effectively, Core-SLS then tries to repair the previous solution to satisfy the new constraints. and (iii) how flipping the variable value would impact the cost of the assignment.
The solution returned by Solve-HS is used to extract the next set of cores. Thus it is important that the solutions returned by Solve-HS are not only of low cost, but guide the IHS solver towards an optimal solution. We incorporate this idea in our SLS integration by ensuring that the solutions are diversified, i.e., that the subsequent solutions returned by Solve-HS have a large Hamming-distance, whenever possible. Hence, we split Core-SLS into two phases: (i) we start from the previous solution and (ii) we restart the search and start from the previous solution but randomly flip a large number of variable values. Whenever we find a solution in both phases, we prefer the one of lower cost and in case of a tie, we prefer the solution from the second phase to foster diversification. Further, Core-SLS’s purpose is avoiding costly calls to the optimizer. Hence, we skip Core-SLS whenever we know from prior iterations that the optimizer calls do not take longer than Core-SLS calls.
0.4 Certified IHS Computations
We now turn our attention to developing proofs for the IHS paradigm, i.e., independently checkable certificates of correctness of the result. The VeriPB format is a general-purpose proof format for pseudo-Boolean optimization, combining implicational reasoning using cutting planes operations (such as adding up two PB constraints) [CCT87] with well-chosen strengthening rules that allow deriving non-implied constraints [BGMN23]. For representational succinctness of the current paper, details of the actual proof rules are not critical to spell out, but it is relevant that the strengthening rules allow us to derive fresh reification variables. In other words, if $x$ is a variable that does not yet appear in the proof and $C$ is a PB constraint, we can always derive (one or both) two new PB constraints $x\Rightarrow C$ and $x\Leftarrow C$ that express that $x$ is true precisely when $C$ is satisfied. If $C$ is the normalized constraint ${\textstyle\sum}_{i}a_{i}\ell_{i}\geq A$, then $x\Rightarrow C$ and $x\Leftarrow C$ are $A\overline{x}+{\textstyle\sum}_{i}a_{i}\ell_{i}\geq A\text{ and }({\textstyle\sum}_{i}a_{i}-A+1)x+{\textstyle\sum}_{i}a_{i}\overline{\ell}_{i}\geq{\textstyle\sum}_{i}a_{i}-A+1\text{, respectively. }$
0.4.1 VeriPB Proofs for PBO-IHS
As instantiations of the conflict-driven pseudo-Boolean approach discussed, e.g., in Section 0.3.1 have had support for VeriPB proof logging for a while (see, e.g., the recent certifying track in the PBO Competitions [Rou24]), a natural idea toward obtaining uniform proofs for PBO-IHS is to instantiate Solve-HS with a PB-based approach and to combine the proofs produced by the Solve-HS subroutine with the proofs produced by the Extract-Cores subroutine. While the idea is conceptually simple, there are a number of details that need to be considered. Due to space restrictions, we cannot go into detail about how the proof production for the different components works. Instead, we discuss challenges in their integration and how they are addressed.
Constraint ID and Variable Management.
With two different solvers co-producing parts of the proof, neither of the solvers are fully aware of which constraints are derived in the proofs and (in case they introduce new variables), which variables are unused so far. To facilitate this, we developed a general pseudo-Boolean proof logging API which both solvers use and which is responsible for maintaining relevant information, such as mappings from solver-specific variables to their name used in the proof. We expect this API to be useful more generally for future pseudo-Boolean proof logging approaches as well.
Non-Reified Solution-Improving Constraints.
Another issue to be dealt with is that the optimizer uses (Solve-HS) is not optimizing the original problem, but a derived problem. When using a pseudo-Boolean optimizer with proof logging, it will log solutions it finds and might then use the associated solution-improving constraint (which expresses that from now on we are only interested in better solutions). Taking a holistic perspective, however, this solution-improving constraint is not sound: the solution Solve-HS finds is not a solution of the whole problem. Above we already described two versions of integrating a SIS solver (incremental and non-incremental). In the incremental case the solver does not introduce a proper solution-improving constraint but a reified version thereof; this is easy to deal with from a proof logging perspective. In the non-incremental version, however, the solver is ran without assumptions and restarted upon every call. In this case, the solver will be deriving those unsound solution-improving constraints. Our solution for dealing with this is to allow for a small discrepancy between the constraints the solver sees and the constraints derived in the proof, making sure the solution-improving constraint logged in the proof is indeed reified.
0.4.2 Proofs and IP Instantiations of Solve-HS
To harness both the effectiveness of SLS and floating-point IP solvers as well as proof-production capabilities of pseudo-Boolean instantiations of Solve-HS, we outline hybrid instantiations of Optimize that combine these strengths. The intuition here is that no proofs are needed for the calls to Solve-HS that do not result in optimal solutions of $(\mathcal{K},O)$. The hybrid schemes we consider seek to minimize the calls to the (less efficient) proof-producing optimizers, while still enabling proofs for the overall IHS algorithm. More specifically, we consider the following hybridizations of non-proof-producing (i.e., SLS or IP) instantiations of Optimize, and the proof-producing ones.
Verify optimal solution (OptLB).
The first approach invokes a proof-producing optimizer whenever the lower bound would be increased to equal the upper bound, i.e., when an optimal solution to the overall instance has been found. Intuitively, this represents the minimum amount of exact computation in Solve-HS needed to obtain a proof of correctness for the overall IHS algorithm; however, no guarantees on the intermediate lower bounds are obtained. Additionally, reinvoking a proof-producing optimizer when the LB would be increased might require some amount of duplicate work in the optimizer recomputing a solution already found by the IP or SLS solver.
Verify all lower bounds (AllLB).
The second approach uses a proof-producing optimizer whenever the lower bound would be refined, i.e whenever the if statement on Line 2 of Algorithm 2 is true. Verifying all of the lower bounds requires more (potentially duplicated) exact computation. Nevertheless, this allows for generating proofs for more search heuristics used by instantiations of IHS that we are aware of.
Switch when forced lower bounds (ForceLB).
The previous two hybridizations possibly perform duplicate work (asking a proof-producing optimizer to confirm all/some optimality claims of other optimizers), which might be undesirable. In this third hybridization, we make an a priori distinction between calls where optimality is needed and calls where this is not needed. More specifically, this version uses proof-producing optimizers to only instantiate the calls to Optimize in which the parameter opt? is true. The calls to Optimize that are not required to find an optimal solution can be instantiated with non-proof producing optimizers. This ensures that any calls to Solve-HS that are required to find an optimal solution are indeed proof-producing. In this setting, the case $\gamma(O)=ub$ of the if statement on Line 2 that represents Optimize returning an optimal solution even when opt? is false, is not used since non-proof-producing optimizers may make such calls.
0.5 Experiments
In our experiments, we evaluate the performance of the different instantiations of Optimize as well as the overhead caused by proof logging. First, we compare how different solvers used as the single solver for Optimize perform as well as the hybrid configurations outlined in Section 0.4. Secondly, we evaluate the benefits of integrating SLS. Thirdly, we investigate the impact of proof logging on runtime performance. All implementation code will be made available in open source, together with run scripts and raw data.
Setup
We run our experiments on compute nodes with two AMD EPYC 7513 CPUs, each with 64 cores running at 2.6 GHz. We reduce runtime variance by reserving whole CPUs for our experiments and run 16 instances in parallel on one CPU. Each run is limited to 14 GB memory and 1 h runtime using runlim [^3] https://github.com/arminbiere/runlim, and each solver is executed once on each instance.
Solvers
We use Gurobi 11.0.3 and SCIP 9.2.3444https://github.com/scipopt/scip (commit 3045f20) [BBB+24] as ILP solvers. The latter offers an exact mode (see Section 0.3.1). Further, we use a modified version of Roundingsat [^5] https://gitlab.com/MIAOresearch/software/roundingsat (Commit: c548e109) [GN21, BGMN23] as the proof producing PB solver and VeriPB 2.2.2666https://gitlab.com/MIAOresearch/software/VeriPB to check these proofs. Proof logging is enabled only for experiments concerning proof logging.
Instances
We evaluate the solvers on the same set of $1786$ instances across tens of different benchmark domains as used in the the original papers presenting PBO-IHS [SBJ21, SBJ22]; see [Smi21] for details.
0.5.1 Impact of Hitting Set Oracles
Hybrid
Single AllLB ForceLB OptLB
Solver # m # m # m # m
Gurobi 951 0.66 - - - - - -
SCIP 913 1.16 - - - - - -
Exact SCIP 762 2.10 - - - - - -
Core Guided 687 0.63 730 0.74 725 0.72 729 0.73
SIS Reified 625 1.00 651 1.00 661 0.98 696 0.81
SIS 672 0.79 684 0.86 697 0.88 696 0.80
SIS + CB 716 2.62 737 1.55 731 2.08 746 1.46
Table 1: Different instantiations of Optimize. Single shows instantiations using only a single solver and Hybrid shows instantations that combine an ILP solver (Gurobi) with a proof producing solver (Roundingsat). #: number of instances solved, m: average memory usage in GB.
In our first experiment, we use a single solver for the hitting set computation in Optimize. Table 1 shows the results for different solvers, split into ILP solvers (top) and designated PB solving methods. These results show that the ILP solvers solver more instances. However, the difference between SCIP and Exact SCIP show that guaranteed optimality causes a significant performance decrease. Indeed, the number of solved instances decreases by more than 15% while the memory usage doubles.
The dedicated PBO methods solve fewer instances, motivating the hybrid approaches. The further we delay the call to RoundingSat, the better the results: AllLB has the most overhead and performs overall poorest, while OptLB requires the fewest RoundingSat calls and performs overall best. Surprisingly, SIS with core boosting does not perform well on its own, being the second-worst method on its own, but performs best in our hybrid approaches. The difference might be due to an overhead in core boosting that only affects performance in frequent incremental calls.
0.5.2 Impact of Integrating Local Search
IHS + SLS OptLB + SLS
Solver # m d # m d
Gurobi +17 -0.02 31 - - -
SCIP +4 -0.14 18 - - -
Exact SCIP +44 -0.48 58 - - -
Core Guided +41 +0.02 53 +11 -0.02 21
SIS Reified +34 -0.12 46 +8 -0.02 30
SIS +13 +0.02 39 +7 -0.02 25
SIS + CB +19 -0.53 53 +3 -0.03 27
Table 2: Different instantiations of Optimize with and without SLS. #: number of solved instances (#); m: average memory usage in GB (positive numbers indicate an increase with SLS); d: number of instances solved with or without SLS but not by both configurations.
We integrated Core-SLS (cf. Section 0.3.2) into the single solver instantiations of Optimize and the best hybrid approach OptLB. By Table 2, SLS improves all instantiations. We use two SLS configurations, one for SLS combined with an ILP solver and one for SLS combined with dedicated PBO approaches. In the hybrid configuration, we use the SLS configuration for ILP. Integration of SLS consistently results in more instances solved, decrease memory usage and improve runtimes. This is most striking for Exact SCIP, for which SLS gives the biggest gain in the number of solved instances, also decreasing the memory consumption by almost 20%. Interestingly, often large numbers of instances are solved either with or without SLS but not both (Column d in Table 2). However, when SLS does not find a sufficiently low-cost solution it directly incurs a runtime overhead. Despite diversification, SLS can get stuck in local optima and cannot escape it without an optimizer. This suggests that improving SLS heuristics further for the IHS use case has potential for even further improvements.
In the hybrids, SLS can only replace ILP solver calls due to inexactness. Here, SLS contributions significantly decrease, as the number of instances that SLS performs well on is balanced with instances it does poorly on. It is unlikely that instances the ILP solver can only solve when SLS is used could be solved by the dedicated PBO approach.
0.5.3 Proof Logging Overhead
We evaluate proof logging using the most performant proof-producing variant, OptLB using SIS + CB. The proofs are checked using VeriPB with a timelimit of 10 h. By proof logging, we found several occasions where the ILP solver wrongly claimed optimality; Gurobi on four instances a total of nine times, and four wrong results for Exact SCIP.
Figure 1: Comparison of runtime with and without proof logging. Times are in seconds.
SIS + CB solved 746 instances without proof logging and 741 with proof logging. The median runtime overhead caused by proof logging is very low, 10.7%, while the overhead at the 90th percentile is 100.0% (see Figure 1). Although beyond the scope of this paper’s contributions, with the present version of the VeriPB proof checker, the median proof checking time is 588.0% of solving with proof logging; 712 proofs could be checked with VeriPB with a 10 h time limit. We expect these numbers to decrease significantly with on-going advances in VeriPB proof checking [Oer25] once extended to optimization.
0.6 Conclusion
The hitting set component in IHS is standardly realized using (typically commercial) IP solvers using floating-point computations. Howver, this can lead to correctness issues in practice due to numerical issues and at least at present hinders gaining trustworthy IHS computations due to difficulties in uniformly certifying the computations of both the core extraction and the hitting set components. With these motivations, focusing on IHS for PBO as one of the most recent successful IHS instantiations, we explored different combinations of recent advances in exact PB reasoning, local search and IP solving for instantiating the hitting set component as potential means of gaining trustworthiness to IHS computations, addressing both issues with numerical instability and the challenge of realizing certified IHS. We extensively evaluate a wide range of different variants of such combinations, and showed that by clever combinations of IP solving, PB reasoning and local search, certified IHS can be realized and uniform proofs obtained for IHS computations while remaining reasonably performant, especially when compared to employing numerically-exact IP solving. The ideas presented are expected to be applicable to various other IHS instantiations, such as state-of-the-art IHS-based solvers for MaxSAT, which is a promising direction for further work.